You purchase 12 shares of Starbucks stock at $67.1 per share. Additionally, you purchase 38 shares of Tyson Foods for $131.7 and 47 shares of Uni-lever for $88.7 each. What is your portfolio weight for Starbucks?

(Enter your response as a percentage with two decimal places, ex: 12.34)
Answer #1

Total value of:

Starbucks stock=(12*67.1)=805.2

Tyson Foods=(38*131.7)=5004.6

Uni-lever=(47*88.7)=4168.9

Total value=(805.2+5004.6+4168.9)=$9978.7

Hence portfolio weight of Starbucks=805.2/9978.7

=8.07%(Approx).